# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v3

@@ Appendix A. Invariant metric on $AdS_{3}$
However, as we see in section two, it is the real form of the Lie algebra $su(1,1)$ or simply the lie algebra $sl(2,R)$ which is in particular relevant for our discussion in relation to the exterior region of the BTZ black hole. Therefore, we obtain below the metric using $sl(2,{\tenrm I\kern-1.79993ptR})$. A convenient choice of basis is $b_{1}=-{1\over 2}\sigma_{1},b_{2}={i\over 2}\sigma_{2},b_{3}=-{1\over 2}\sigma_{3}$. The analogous parametrization is
@@ Appendix B. Bosonic T-duality
where $\widehat{m}$ takes all values of $m$ except $m=1$. Integrating out the field $\chi$, i.e., eliminating it using its equation of motion, gives back (B.1). In general a one form on $S^{1}$ is not exact. However, since $X^{1}$ is compact and thus it winds around in either directions integer times, we can still write the gauge potential on $S^{1}$ as the gradient of a compact scalar and make the above identifications. This will get us back to the original action.
@@ Appendix C. The Brown-York quasi-local energy and mass
In this appendix we review the Brown-York quasi-local energy and mass [5]. We primarily follow the work [5], but see also [4] for a related insightful discussion. We specialize the discussion to $(2+1)$ dimensions. We furthermore restrict ourselves to axially symmetric stationary solutions. Therefore, we can directly use the formulae in section four of the paper.
@@
where $A,B,C$ and $Z$ are functions of only the radial coordinate $r$. The coordinate $\theta$ is an angle variable and it has $2\pi$ periodicity. The spacetime contains no inner boundaries. It can have event horizons [4].21 In the case the spacetime contains other null boundaries we include appropriate terms in the action functional that contribute on these boundaries. This will not concern us in this paper. We refer to [46] and references therein for such cases. We assume the metric is minimally coupled to a scalar. Therefore, the scalar action contains no derivatives of the metric. We also assume the scalar depends only on $r$.
Let $u^{\mu}$ denote the future pointing time-like unit normal to the $t={\tenrm constant}$ two-space denoted by $\Sigma_{t}$. $t$ labels the two-spaces $\Sigma_{t}$ and it defines the time direction or coordinate. This particular slicing is the usual ADM slicing of spacetime [47]. We denote the (intrinsic) metric on $\Sigma_{t}$ by $h_{ij}$.22 For convenience, here and below we use lower latin letters $i=t,r,\theta$ to denote the intrinsic metrics and we use greek letters $\mu=t,r,\theta$ to denote the induced metrics. We denote the boundary of $\Sigma_{t}$ by $B_{t}$. The world-line or history of $B_{t}$ is denoted ${}^{2}B$. Let $n^{\mu}$ is the outward pointing space-like unit normal to the two-boundary ${}^{2}B$. For simplicity, we assume on the time-like boundary ${}^{2}B$ that $n^{\mu}u_{\mu}|_{{}^{2}B}=0$. That is, the surface $\Sigma_{t}$ meets the boundary ${}^{2}B$ orthogonally. This is not however essential or necessary for the analysis below [[48]48]. This is in part because we are ultimately interested in the case ${}^{2}B$ is at radial infinity where the spacetime (C.1) is (conformally) flat. We denote the metric on ${}^{2}B$ by $\gamma_{ij}$. We denote the metric on $B_{t}$ which is the intersection of ${}^{2}B$ and $\Sigma_{t}$ by $\sigma_{ij}$.
@@
$\eqalign{h_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu}\cr\gamma_{\mu\nu}&=g_{\mu\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu},\cr\sigma_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu}.}$ (C.3)C.3( C. 3 )
@@
We now give the conserved charges in terms of the components of the metric. The energy associated with the Killing field $\xi^{i}=u^{i}=\gamma_{i}^{\mu}u_{\mu}$23 Note ${\cal D}_{(i}\xi_{j)}=\gamma_{(i}^{\lambda}\nabla_{\lambda}\gamma_{j)}^{\mu}u_{\mu}=0$. is given by
@@
In gravity theories physical quantities such as energy are defined only with respect to a reference. Thus, we must specify a reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$, where $g_{0}$ denotes the metric and $\phi_{0}$ collectively denotes the matter fields on ${\cal M}_{0}$. $\Lambda$ is possible cosmological constant. The reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is a particular solution to Einstein equations. It is chosen such that the deviations
@@
vanish smoothly at the boundary $B_{t}$ [3]. In the case ${\cal M}(g,\phi;\Lambda)$ is asymptotically flat spacetime, the reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is usually taken to be the Minkowski spacetime. In the paper in all the cases we study the reference metrics are conformally flat spacetimes.
@@
We will assume $A_{0},B_{0},C_{0}$ and $Z_{0}$ are functions of only $r$. The scalar field depends only on $r$ and it is coupled to the metric minimally.
The reference metrics $g_{0}$ that we consider in this paper have $Z_{0}={\tenrm constant}$. Accordingly, the (quasi-local) energy is
@@
The (quasi-local) mass is
@@
In section four of the paper we set $\kappa^{2}=\pi$ for simplicity. This is equivalent to $18G_{N}=$ 1.
@@ Appendix D. Kruskal coordinates
The metrics which we study in this paper are in general of the form $ds^{2}=\Delta^{2}d{\widehat{s}}^{2}$, where the factor $\Delta^{2}$ is (manifestly) positive and regular everywhere.25 Except at a curvature singularity. The line element $d{\widehat{s}}^{2}$ is of two types. One type has the form (see, e.g., (3.15) and (4.14))
@@
$\rho^{\star}={1\over 2\kappa}\ln\left({u^{2}-v^{2}\over 4A^{2}}\right),\quad t={1\over 2\kappa}\tanh^{-1}\left({2uv\over u^{2}+v^{2}}\right)$ (D.13)D.13( D. 13 )
@@
In the large $x$ limit we get $x^{\star}=2|c_{2}|c_{1}^{2}\ln x+h(c_{1},c_{2},\lambda)$.26 The BTZ solution is obtained by setting $\mu=1$, i.e., $c_{2}=0$. For $c_{2}=0$, in the large $x$ limit, we instead obatin $x^{\star}=-c_{1}^{2}/x+{\cal O}(x^{-3})$. The function $h$ in particular satisfies $h(c_{1},0,\lambda)=0$. In the large $x$ limit and at constant $\varphi$ the metric (D.3) up to an overall positive factor equals $du^{2}-dv^{2}+[\lambda^{2}/(u^{2}-v^{2})](udv-vdu)^{2}$. Therefore, outgoing and ingoing radial null lines end on future and past null infinities, respectively.
@@
Fig. 7: The plots depict the Penrose diagrams of the regions I, II and III. Each point represents a circle or the orbit of the isometry $\partial_{\varphi}$. Straight lines at $\pi/4$ radians represent null lines. We used the values $\lambda=0.4,\mu=0.5,\rho_{+}=1,l=1,A=1$. For $A=-1$ we simply flips $U\to-U$. Plot a) is the Penrose diagram for III. $x=0$ is mapped to the vertical line $U=\pi/2$. Plot b) is the Penrose diagram for II. It is obtained using $K_{2}$. The lines $u=\pm v$ at $x=\lambda$ are mapped to the lines $U=\pm V$. $x=1$ is mapped to $U\pm V=\pi$. Plot c) is the Penrose diagram for I. Null lines at $x=1$ are mapped to $U=\pm V$. $x=\infty$ is mapped to $U\pm V=\pi$.

### v4

@@ Appendix A. Invariant metric on $AdS_{3}$
However, as we see in section two, it is the real form of the Lie algebra $su(1,1)$ (i.e., now the field is ${\tenrm I\kern-1.79993ptR}$ however the basis can be complex) or simply the lie algebra $sl(2,R)$ which is in particular relevant for our discussion in relation to the exterior region of the BTZ black hole. Therefore, we obtain below the metric using $sl(2,{\tenrm I\kern-1.79993ptR})$. A convenient choice of basis is $b_{1}=-{1\over 2}\sigma_{1},b_{2}={i\over 2}\sigma_{2},b_{3}=-{1\over 2}\sigma_{3}$. The analogous parametrization is
@@
We simply replace $\theta$ with $\overline{\theta}\in{\tenrm I\kern-1.79993ptR}$.
@@ Appendix B. Bosonic T-duality
where $\widehat{m}$ takes all values of $m$ except $m=1$. Integrating out the field $\chi$, i.e., eliminating it using its equation of motion, gives back (B.1). In general a one form on $S^{1}$ is not exact. However, since $X^{1}$ is compact and thus it winds around in either directions integer times, we can still write the gauge potential on $S^{1}$ as the gradient of a compact scalar $A=\partial\chi^{1},{\overline{A}}={\overline{\partial}}\chi^{1}$ and make the identification $\chi^{1}\sim X^{1}$. This will get us back to the original action.
@@ Appendix C. The Brown-York quasi-local energy and mass
In this appendix we review the Brown-York quasi-local energy and mass [5]. We primarily follow the work [5], but see also [4] for a related insightful discussion. We specialize the discussion for definiteness to $(2+1)$ dimensions. We furthermore restrict ourselves to axially symmetric stationary solutions. Therefore, we can directly use the formulae in section four of the paper.
@@
where $A,B,C$ and $Z$ are functions of only the radial coordinate $r$. The coordinate $\theta$ is an angle variable and it has $2\pi$ periodicity. The spacetime contains no inner boundaries. It can have event horizons [4].22 In the case the spacetime contains other null boundaries we include appropriate terms in the action functional that contribute on these boundaries. This will not concern us in this paper. We refer to [46] and references therein for such cases. We assume the metric is minimally coupled to a scalar. Therefore, the scalar action contains no derivatives of the metric. We also assume the scalar depends only on $r$.
Let $u^{\mu}$ denote the future pointing time-like unit normal to the $t={\tenrm constant}$ two-space denoted by $\Sigma_{t}$. $t$ labels the two-spaces $\Sigma_{t}$ and it defines the time direction or coordinate. This particular slicing is the usual ADM slicing of spacetime [47]. We denote the (intrinsic) metric on $\Sigma_{t}$ by $h_{ij}$.23 For convenience, here and below we use lower latin letters $i=t,r,\theta$ to denote the intrinsic metrics and we use greek letters $\mu=t,r,\theta$ to denote the induced metrics. We denote the boundary of $\Sigma_{t}$ by $B_{t}$. The world-line or history of $B_{t}$ is denoted ${}^{2}B$. Let $n^{\mu}$ is the outward pointing space-like unit normal to the two-boundary ${}^{2}B$. For simplicity, we assume on the time-like boundary ${}^{2}B$ that $n^{\mu}u_{\mu}|_{{}^{2}B}=0$. That is, the surface $\Sigma_{t}$ meets the boundary ${}^{2}B$ orthogonally. This is not however essential or necessary for the analysis below [[48]48]. This is in part because we are ultimately interested in the case ${}^{2}B$ is at radial infinity where the spacetime (C.1) is (conformally) flat. We denote the metric on ${}^{2}B$ by $\gamma_{ij}$. We denote the metric on $B_{t}$ which is the intersection of ${}^{2}B$ and $\Sigma_{t}$ by $\sigma_{ij}$.
@@
$\eqalign{h_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu},\cr\gamma_{\mu\nu}&=g_{\mu\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu},\cr\sigma_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu}.}$ (C.3)C.3( C. 3 )
@@
We now give the conserved charges in terms of the components of the metric. The energy associated with the Killing field $\xi^{i}=u^{i}=\gamma_{i}^{\mu}u_{\mu}$24 Note ${\cal D}_{(i}\xi_{j)}=\gamma_{(i}^{\lambda}\nabla_{\lambda}\gamma_{j)}^{\mu}u_{\mu}=0$. is given by
@@
In gravity theories physical quantities such as energy are defined only with respect to a reference. Thus, we must specify a reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$, where $g_{0}$ denotes the metric and $\phi_{0}$ collectively denotes the matter fields on ${\cal M}_{0}$. $\Lambda$ is possible cosmological constant.
The reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is a particular solution to Einstein equations. It is chosen such that the deviations or fluctuations
@@
vanish smoothly at the boundary $B_{t}$ [3]. Note $B_{t}$ is the boundary of the $t={\tenrm constant}$ spacetime slice. In the case ${\cal M}(g,\phi;\Lambda)$ is asymptotically flat spacetime, the reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is usually taken to be the Minkowski spacetime. In this paper in all the cases we study the reference metrics are conformally flat spacetimes.
@@
We will assume $A_{0},B_{0},C_{0}$ and $Z_{0}$ are functions of only $r$. The scalar field depends only on $r$ and it is coupled to the metric minimally. At the boundary $B_{t}$, $h_{\mu\nu}=0$ and $\chi=0$ (C.23). That is, $A(r^{\star})=A_{0}(r^{\star}),B(r^{\star})=B_{0}(r^{\star}),C(r^{\star})=C_{0}(r^{\star}),Z(r^{\star})=Z_{0}(r^{\star})$ and $\phi(r^{\star})=\phi_{0}(r^{\star})$, where $r^{\star}$ is the location of $B_{t}$.
The reference metrics $g_{0}$ that we consider in this paper have $Z_{0}={\tenrm constant}$. Accordingly, since $\partial_{r}Z_{0}=0$, the (quasi-local) angular momentum remains unchanged. It is still given by (C.21). The (quasi-local) energy is given by26 In general it is quasi-local in the sense that it is local only on the (hyper-)surface $\Sigma_{t}$, i.e., the $t={\tenrm constant}$ spacetime slice. It follows, in $(2+1)$ dimensions, it is defined as a local quantity.
@@
We implicitly assume the limit $r\to r^{\star}$ exists. The (quasi-local) mass is
@@
In section four of the paper we set $\kappa^{2}=\pi$ (in some length unit) for simplicity. This is equivalent to $18G_{N}=$ 1 (in some length unit).
@@ Appendix D. Kruskal coordinates
The metrics which we study in this paper are in general of the form $ds^{2}=\Delta^{2}d{\widehat{s}}^{2}$, where the factor $\Delta^{2}$ is (manifestly) positive and regular everywhere.27 Except at a curvature singularity. The line element $d{\widehat{s}}^{2}$ is of two types. One type has the form (see, e.g., (3.15) and (4.14))
@@
$\rho^{\star}={1\over 2\kappa}\ln\left({u^{2}-v^{2}\over 4A^{2}}\right),\quad t={1\over 2\kappa}\tanh^{-1}\left({2uv\over u^{2}+v^{2}}\right).$ (D.13)D.13( D. 13 )
@@
In the large $x$ limit we get $x^{\star}=2|c_{2}|c_{1}^{2}\ln x+h(c_{1},c_{2},\lambda)$.28 The BTZ solution is obtained by setting $\mu=1$, i.e., $c_{2}=0$. For $c_{2}=0$, in the large $x$ limit, we instead obatin $x^{\star}=-c_{1}^{2}/x+{\cal O}(x^{-3})$. The function $h$ in particular satisfies $h(c_{1},0,\lambda)=0$. In the large $x$ limit and at constant $\varphi$ the metric (D.3) up to an overall positive factor equals $du^{2}-dv^{2}+[\lambda^{2}/(u^{2}-v^{2})](udv-vdu)^{2}$. Therefore, outgoing and ingoing radial null lines end on future and past null infinities, respectively.
@@
Fig. 7: The plots depict the Penrose diagrams of the regions I, II and III. Each point represents a circle or the orbit of the isometry $\partial_{\varphi}$. Straight lines at $\pi/4$ radians represent null lines. We used the values $\lambda=0.4,\mu=0.5,A=1$, $\rho_{+}=1$ (in string units), $l=1$ (in string units). For $A=-1$ we simply flips $U\to-U$. Plot a) is the Penrose diagram for III. $x=0$ is mapped to the vertical line $U=\pi/2$. Plot b) is the Penrose diagram for II. It is obtained using $K_{2}$. The lines $u=\pm v$ at $x=\lambda$ are mapped to the lines $U=\pm V$. $x=1$ is mapped to $U\pm V=\pi$. Plot c) is the Penrose diagram for I. Null lines at $x=1$ are mapped to $U=\pm V$. $x=\infty$ is mapped to $U\pm V=\pi$.